# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 8.1 Explanation-guided Data Augmentation for Mitigating Shortcuts
Machine learning models are prone to make predictions with spurious correlations, also known as shortcuts (Geirhos et al., 2020), which are misaligned with human reasoning processes. This dependency on shortcuts underlies various challenges in machine learning, notably diminishing a model’s ability to generalize to out-of-distribution samples and weakening its resilience against adversarial attacks (Xu et al., 2020). The extent to which a model depends on shortcut features is intrinsically linked to its interpretability. Essentially, if a model’s predictions are predominantly based on such unreliable features, it indicates that the underlying mechanisms of the model are not fully comprehensible or interpretable from a human perspective. Therefore, post-hoc explanation techniques usually play a crucial role in detecting undesirable correlations between input and predictions within deep models (Liu et al., 2018; 2021). For example, Du et al. (2021) adopt Integrated Gradient (IG) to attribute a model’s predictions to its input features, showing that the model tends to treat functional words, numbers, and negation words as shortcuts and strongly rely on these shortcut features for prediction in natural language understanding tasks.
Data augmentation can be applied to training better models that are less sensitive to shortcut features. Explanatory information such as counterfactuals (Wang & Culotta, 2021) has been incorporated in data augmentation to improve model robustness. It generates counterfactual samples by first identifying critical features (e.g., word tokens) and then replacing these features with their antonyms, along with reversing their associated labels. Subsequently, the generated samples are combined with the original ones to train downstream models. Furthermore, these techniques can be extended to enhance the out-of-distribution performance of smaller models (Sachdeva et al., 2023; Wen et al., 2022). Namely, large language models could serve as an effective tool to augment data. For example, LLMs are able to synthesize examples that represent outlier cases or rare occurrences, which helps smaller models generalize better on unseen data (Xu et al., 2023a). This could be helpful in building robust models in scenarios where data are scarce or confidential (Tang et al., 2023a). Besides, LLMs are promising in improving models’ safety by generating adversarial examples that are more valid and natural compared to conventional approaches (Wang et al., 2023f). First, the most vulnerable words are identified with attribution-based methods. Then, these words are replaced by LLMs in a manner that maintains the integrity of the original text. The quality of these examples can be examined with an external classifier. Subsequently, these adversarial examples are employed to train downstream models, effectively fortifying them against potential attacks and boosting their security. Similarly, LLMs are also helpful in mitigating biases such as fairness issues in models (He et al., 2023). This work claims that it automatically identifies underrepresented subgroups, and chooses an optimal augmentation strategy that avoids hurting other groups. New group examples are generated using LLMs with human-providing labels. The experiments observe improvements on both underrepresented groups and overall model performance. This approach prevents skewed datasets from leading to unfair outcomes for a certain demographic, thereby potentially promoting fairness in society.
8.2 Explanation-enhanced Data Enrichment
As a powerful generative tool, LLMs have been leveraged to directly generate natural language explanations as augmented data. This strategy relies on LLMs’ understanding abilities to assist smaller models in their learning tasks. One objective in such work is to add natural language explanation generated by LLMs to training data, so as to enhance the performance of small models. Li et al. (2022b) introduce explanations from LLMs that facilitate the training of smaller models to enhance their reasoning capabilities and acquire explanation generation abilities. Specifically, three explanation generation approaches from LLM are utilized to train smaller models, including (1) explanations generated through chain of thought prompting, (2) explanations by rationalization prompting conditioned on golden labels and (3) hybrid approach that combines the first two approaches. With the best case scenario in a reasoning case, a 9.5% improvement on accuracy compared to 60x larger GPT-3 has been seen (Li et al., 2022b). It is worth noting that LLMs including ChatGPT and GPT-4 are competitive in generating coherent and logical explanations, except for limited languages (Whitehouse et al., 2023). Explanations from LLMs have also been utilized to mitigate spurious correlations in aspect-based sentiment analysis tasks (Wang et al., 2023c). This study proposes using LLMs to generate explanations for the sentiment labels of aspects in sentences (Wang et al., 2023c). These explanations provide reasoning grounded in contextual semantics rather than relying on spurious correlations between words and labels. The explanations are integrated into the training of aspect-based sentiment analysis models through two methods: augmenting the training data with the explanations or distilling knowledge from the explanations through mimicking behaviors. By focusing on the explanations instead of superficial textual cues, the models can better learn the true associations between text and sentiment and become more robust, improving both in-domain performance and generalization ability (Wang et al., 2023c). Another line of work involves integrating LLM rationales as additional supervision to guide the training of smaller models. Experiments have shown that this approach not only requires fewer training data but also outperforms traditional fine-tuning and distillation methods (Hsieh et al., 2023).
In addition to the existing application of augmentation techniques summarized above, we envision that incorporating additional text information can also be practical and efficient in enhancing the performance of various models. For example, one promising application lies in the realm of guiding the parameter learning process of small models by using automatically generated explanations. Previous research has investigated this avenue by directing the attention of natural language inference models towards human-crafted explanations (Stacey et al., 2022). As human-crafted explanations are both arduous and non-transferable, utilizing LLMs as generators presents a more economic and versatile alternation. Another potential application is to enhance model performance on complex tasks using natural language explanations from LLMs. For instance, code translation generation tasks incorporate explanations as an intermediate step, improving model performance by 12% on average (Tang et al., 2023b). The result shows that explanations are particularly useful in zero-shot settings. Apart from assisting in smaller models, LLMs have demonstrated their ability to improve their own reasoning abilities by generating reliable rationales (Huang et al., 2022). Further, Krishna et al. (2023) embed post-hoc explanations, attributing scores to all input features, into natural language rationales. This approach improves model accuracy by 10-25% across multiple tasks via in-context learning. Another study explicitly investigates LLMs’ ability to generate post-hoc explanations in natural language. The experiments show that LLMs spot key features with 75% accuracy (Kroeger et al., 2023). These studies present a novel strategy for leveraging post-hoc explanations, beyond traditional natural language explanations, to enrich training data so as to bolster model performance.
8.3 Challenges
8.3.1 Computational Overhead
@@
Conventional post-hoc explanations, built on well-trained models, are often resource-intensive tasks. The first scenario mentioned above leverages interpretability techniques to accurately diagnose dataset issues. This process typically requires multiple rounds of model training and applying interpretability methods to develop fair and robust models. Consequently, the crafting process can be both time and energy-consuming. Given these challenges, exploring the development of data-centric evaluation metrics is crucial. These metrics can offer a more efficient way to assess data issues, bypassing traditional, cumbersome explanation methods. By focusing on these data-centric measurements, data issues can be diagnosed and fixed before training. The number of training rounds needed is then significantly reduced. This shift not only streamlines model development but also helps reduce computational overhead, making the whole process more practical and efficient.
8.3.2 Data Quality and Volume
@@
Despite their advanced capabilities, LLMs still have limitations when dealing with highly specialized or niche contexts. For example, one of the most prominent issues is “hallucination”, where models generate plausible but incorrect or misleading responses. This could adversely affect the quality of augmented data, potentially introducing more biases to which LLMs are also vulnerable. Another challenge is controlling the relevance of LLM-generated content. That is, the explanations or data points may seem reasonable but often lack factual accuracy or nuances specific to a domain. Currently, we lack robust metrics to effectively measure the quality and relevance of these generated data relative to the original tasks. Determining the precise amount of data required is also challenging, often leading to new dataset imbalances. Managing the quality of LLM-generated data is an immense challenge, as augmented data can introduce other biases (Zhao & Vydiswaran, 2021). This stems from LLMs’ limited ability to accurately control the quantity and distribution of generated data. Moreover, crafting effective prompts is more of an art than a science, adding uncertainty around generated data quality. Together, these factors underscore the complexities and challenges in fully harnessing LLMs’ potential for data augmentation and related tasks.
9 Generating User-Friendly Explanation for XAI
@@
Previous sections mainly focused on quantitative explanations with LLM via numerical values. For example, sample-based explanation discussed in Section 4 aims to assign each training sample an influence score (see Eqs.1-4) that measures the confidence that we can use that training sample to explain the prediction of a test sample. However, using numerical values for explanations is not intuitive, which can be difficult to understand by practitioners with little domain knowledge (Latif & Zhai, 2024; Lee et al., 2023; Li et al., 2020). User-friendly explanations, on the contrary, aim to generate human-understandable explanations, e.g., natural language-based descriptions, regarding certain data, the reason why a model makes certain predictions, or what role a neuron plays in the network, such that the generated explanations can be well-understood by both researchers and practitioners.
Given an explainee $e$, which can be a data sample $(x_{i},y_{i})$, a neuron $\theta_{i}$ from a pretrained model $f_{\theta}$, or a prediction result $\hat{y}$ based on the input $x$, generating user-friendly explanation aims to map the explainee $e$ to a sequence of natural language tokens as the explanation for the explainee $e$, such that the generated explanations can be easily comprehended by human beings.
9.1 User-friendly Data Explanation with LLMs
Data explanation refers to the process of translating difficult materials (e.g., program codes, long documents) into concise and straightforward language so that they are easy to understand by humans. Language models have long been used to generate explanations for textual data (Dai & Callan, 2019). Since modern LLMs are trained on corpora composed of codes, math, and papers, they can be leveraged to explain data beyond pure textual content. For example, Chen et al. (2021) have demonstrated that pretrained GPT models possess the ability to understand and generate codes, where explanatory comments are generated simultaneously that facilitate the understanding of programmers. In addition, Welleck et al. (2022) propose to explain math theorems by providing detailed derivations, so that the theorems are easier to understand. Recently, LLMs have also been used to elucidate academic papers (Castillo-González et al., 2022), making difficult content to be easily understood by individuals with little domain knowledge.
9.2 Explaining Small Models with LLMs
@@
Recently, there has been growing interest in leveraging LLMs to generate free-text explanations for small models. For example, to explain black-box text classifiers, Bhattacharjee et al. (2023) propose a prompting-based strategy to identify keywords $K=\{k_{1},k_{2},...,k_{n}\}$ in the input texts $x$ with pretrained LLMs that are informative for the label $y$, and ask LLMs to substitute them with another set of keywords $K^{\prime}=\{k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{n}\}$, such that changed text $x^{\prime}$ changes the label prediction to $y^{\prime}$. They view the textual mapping rule “if we change $K$ into $K^{\prime}$ in $x$, then $y$ will be classified as $y^{\prime}$” as the counterfactual explanation for the model. In addition, to explain the neuron of a pretrained language model (e.g., GPT2), Bills et al. (2023) propose to summarize the neuron activation patterns into textual phrases with a larger language model (e.g., GPT4), where the neuron activation patterns are expressed as a sequence of (token, attribution score) pairs. To verify the identified patterns, they generate activation patterns according to the phrases via the same LLM and compare their similarity with the true activation patterns of the neuron, where the phrases with high scores are considered more confident to serve as the explanation for the neuron.
@@
The explaining ability of LLMs is not necessarily limited to text models. For example, Zhao et al. (2023a) propose using pretrained vision-language models to generate explanations for a neuron $\theta_{i}$ of an image classification model. Specifically, for each class $y=y_{c}$, they first find regions in images with label $y_{c}$ that have maximum activation of the neuron $\theta_{i}$ as the surrogate explainees for $\theta_{i}$, and prompt LLMs such as ChatGPT to generate candidate explanations (words, short phrases) for the class label $y_{c}$. Then, they use the pretrained vision-language model CLIP (Radford et al., 2021) to match the candidate explanations with the surrogate explainees as the explanations for the neuron $\theta_{i}$. Recently, LLMs have also found applications in explaining recommender systems Zhu et al. (2023a). Specifically, Yang et al. (2023c) found that LLMs can well interpret the latent space of sequential recommendation model after alignment, whereas Lei et al. (2023) propose to align user tokens of LLMs with the learned user embeddings of small recommendation model to generate explanations of user preferences encoded in the embeddings. Recently,  Schwettmann et al. (2024) propose a unified framework to explain all models where inputs and outputs can be converted to textual strings. Specifically, the explainer LLM is used as an agent to interact with the explainee model by iteratively creating inputs and observing outputs from the model, where the textual explanations are generated by viewing all the interactions as the context.
9.3 Self-Explanation of LLMs
@@
Due to the black-box nature of LLMs, it is promising to generate user-friendly explanations for the LLMs themselves, such that the operational mechanics and the predictions of LLMs can be well-understood by human experts. Based on whether the LLM needs to be retrained to generate explanations for themselves, the self-explanation of LLM can be categorized into two classes: fine-tuning based approach and in-context based approach, which will be introduced in the following parts.
@@
Fine-tuning based approaches. Given sufficient exemplar explanations on the labels of the training data (e.g., in recommendation datasets such as the Amazon Review datasets (He et al., 2017) or the Yelp dataset (Zhou et al., 2020), users have provided explanations on why they have purchased certain items, which can be viewed as explanations for the ratings), LLMs can learn to generate explanations for their predictions as an auxiliary task through supervised learning. One exemplar method is P5 (Geng et al., 2022), which fine-tunes the pre-trained language model T5 (Raffel et al., 2020) on both the rating and explanation data to generate an explanation alongside the recommendations. Recently, several works have improved upon P5 (Cui et al., 2022; Zhu et al., 2024), which fine-tunes different LLMs such as GPT2, LLaMA, Vicuna, etc., and propose different prompt learning strategies (Li et al., 2023d) with generating explanation as the auxiliary task. With explanations introduced as additional supervision signals to fine-tune pretrained LLMs for recommendations, the performance can be improved with good explainability.
@@
In-context based approaches. In many applications, there is often a lack of sufficient exemplar explanations. However, the unique capability of modern LLMs to reason and provide answers through human-like prompts introduces the potential for in-context based explanations. Here, explanations for predictions are crafted solely based on the information within the prompt. A leading approach in this domain is the Chain-of-Thoughts (CoT) prompting (Wei et al., 2022), which provides few-shot examples (with or without explanations) in the prompt and asks the LLM to generate answers after reasoning step-by-step, where the intermediate reasoning steps that provide more context for generating the final answer can be viewed as explanations. However, CoT generates reasoning first and then based on which generates predictions, where the reasoning steps can influence prediction results (Lyu et al., 2023). If explanations are generated after the prediction, since the explanation is conditioned on the predicted label, it can provide a more faithful post-hoc explanation of why the model makes certain decisions (Lanham et al., 2023). The application of in-context based self-explanation of LLMs is broad. For example, Huang et al. (2023d) explore generating zero-shot self-explanation of sentiment analysis with LLMs by directly asking them to generate explanations alongside the predictions. In addition, Huang et al. (2023a) propose a chain-of-explanation strategy that aims to explain how LLMs can detect hate speech from the textual input. Lu et al. (2022) find that CoT can generate well-supported explanations for question answering with scientific knowledge.
9.4 Challenges
9.4.1 Usability v.s. Reliability

### v2

@@ 8.1. Label-based Data Augmentation
In traditional classification tasks, each data point $x_{i}$ has a label $y_{i}$ representing a class. Models are trained to predict the labels for new data points. However, these models are prone to make predictions with spurious correlations also known as shortcuts (Geirhos et al., 2020). Thus, model’s predictions are predominantly based on shortcut features, and the underlying mechanisms of the model are not interpretable from a human perspective. Post-hoc explanation techniques play a crucial role in detecting undesirable correlations between input and predictions (Liu et al., 2018, 2021). For example, Du et al. (2021) adopt Integrated Gradient (IG) to attribute a model’s predictions to its input features, showing that the model tends to treat functional words, numbers, and negation words as shortcuts and rely on them for prediction in natural language understanding tasks.
LLMs can be applied to resolve data set issues by augmenting various data. For example, explanatory information such as counterfactuals has been incorporated to improve model robustness (Wang and Culotta, 2021) and out-of-distribution performance (Sachdeva et al., 2023). LLMs can also synthesize rare examples, which helps models to generalize better on unseen data (Xu et al., 2023b). Besides, another work utilizes attribution-based interpretability tool to identify vulnerable words and leverage LLMs to synthesize natural adversarial examples to enhance model safety (Wang et al., 2023c). Similarly, LLMs are also helpful in mitigating biases such as fairness issues in models. He et al. (2023) claims that it automatically identifies underrepresented subgroups, and use LLMs to augment theese subgroups while avoids hurting other groups.
In text generation tasks, each data point consists of a prompt $x$ and a response $y$. The data used for model training is usually text including both prompt and response. Generative models are designed to predict next token based on existing input, which can be formulated as $y_{m}:\arg\max p_{\theta}(y_{m}|x,y_{(i\ :\ m-1)})$. To make models generate desired output, well-aligned models are indispensable. Recently, a few work employ LLMs to augment data and enhance model alignment. For example, Wang et al. (Wang et al., 2024b) identifies the weakness of existing datasets and instructs LLMs to generate desired data to improve model safety. Another work use attribution-based method to identify key information from privacy-relevant data and replace the privacy information with synthetic information, which provides a way to avoid privacy leakage (Zeng et al., 2024b).
8.2. Rationale-based Data Enrichment
LLMs have been leveraged to directly generate natural language explanations as augmented data. This strategy makes use of LLMs’ chain-of-thought reasoning abilities (Kojima et al., 2022). The training datasets contain data with annotated information, denotes as $\mathcal{D}=\{(x_{1},r_{1},y_{1}),\ \cdots,(x_{m},r_{m},y_{m})\}$. Each data point includes not only prompt $x$ and response $y$, but reasoning process or annotated information defined as $r$. As state-of-art LLMs are excellent in interpreting context, many studies have utilized them to provide annotations and enrich data.
One line of research uses generated explanations to assist and guide model training.  Li et al. (2022a) introduce LLM explanations to facilitate the training of smaller models, which improves their reasoning capabilities and acquires explanation generation abilities. Hsieh et al (Hsieh et al., 2023) integrate LLM rationales as additional supervision to guide the training of smaller models. Experiments have shown that this approach not only requires fewer training data but also outperforms traditional fine-tuning and distillation methods. Another line of study employ annotated information to boost models’ performance on specific tasks. For example, LLM explanations have also been utilized to generate explanations for sentiment labels of aspects in sentences to mitigate spurious correlations in aspect-based sentiment analysis tasks (Wang et al., 2023a). Besides, Code translation generation tasks incorporate explanations as an intermediate step, improving model performance by 12% on average (Tang et al., 2023). The result shows that explanations are particularly useful in zero-shot settings. Wei et al. (Wei et al., 2024a) instruct LLMs to explain the retrieval process from documents to produce rationales. Then they can be used to train models and further improve models ability on retrieval-augmented generation tasks. Similarly, Lupidi et al. (Lupidi et al., 2024) prompt LLMs to generate data explanations from real-world sources such as articles, which are then used to fine-tune models and improve their performance on multi-hop reasoning and information retrieval.
8.3. Challenges
8.3.1. Computational Overhead
@@
Conventional post-hoc explanations, built on well-trained models, are often resource-intensive tasks. The first scenario mentioned above leverages interpretability techniques to accurately diagnose dataset issues. This process typically requires multiple rounds of model training and applying interpretability methods to develop fair and robust models. Consequently, the crafting process can be both time- and energy-consuming.
8.3.2. Data Quality and Volume
@@
Despite their advanced capabilities, LLMs still have limitations when dealing with highly specialized or niche contexts. For example, one of the most prominent issues is “hallucination”, where models generate plausible but incorrect or misleading responses. This could adversely affect the quality of augmented data, potentially introducing more biases to which LLMs are also vulnerable. Determining the precise amount of data required is also challenging, often leading to new dataset imbalances. Managing the quantity of LLM-generated data is an immense challenge, as augmented data can introduce other biases (Zhao and Vydiswaran, 2021). This stems from LLMs’ limited ability to accurately control the quantity and distribution of generated data. Together, these factors underscore the complexities and challenges in fully harnessing LLMs’ potential for data augmentation and related tasks.
9. Generating User-Friendly Explanation with LLMs
@@
Previous sections mainly focused on quantitative explanations with LLM via numerical values. For example, sample-based explanation discussed in Section 4 aims to assign each training sample an influence score (see Eqs.8-11) that measures the confidence that we can use that training sample to explain the prediction of a test sample. However, using numerical values for explanations is not intuitive, which can be difficult to understand by practitioners with little domain knowledge (Latif and Zhai, 2024; Lee et al., 2023; Li et al., 2020; Chen et al., 2024c). User-friendly explanations, on the contrary, aim to generate human-understandable explanations, e.g., natural language-based descriptions, regarding why a model makes certain predictions or what role a neuron plays in the network, such that the explanations can be well-understood by both researchers and practitioners.
9.1. Explaining Small Models with LLMs
@@
Recently, there has been growing interest in leveraging LLMs to generate free-text explanations for small models. For example, to explain black-box text classifiers, (Bhattacharjee et al., 2023) propose a prompting-based strategy to identify keywords $K=\{k_{1},k_{2},...,k_{n}\}$ in the input texts $x$ with pretrained LLMs that are informative for the label $y$, and ask LLMs to substitute them with another set of keywords $K^{\prime}=\{k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{n}\}$, such that changed text $x^{\prime}$ changes the label prediction to $y^{\prime}$. They view the textual mapping rule “if we change $K$ into $K^{\prime}$ in $x$, then $y$ will be classified as $y^{\prime}$” as the counterfactual explanation for the model. In addition, to explain the neuron of a pretrained language model (e.g., GPT2), Bills et al. (2023) propose to summarize the neuron activation patterns into textual phrases with a larger language model (e.g., GPT4), where the neuron activation patterns are expressed as a sequence of (token, attribution score) pairs. To verify the identified patterns, they generate activation patterns according to the phrases via the same LLM and compare their similarity with the true activation patterns of the neuron, where the phrases with high scores are considered more confident to serve as the explanation for the neuron.
@@
The explaining ability of LLMs is not necessarily limited to text models. For example, (Zhao et al., 2023b) propose using pretrained vision-language models to generate explanations for a neuron $\theta_{i}$ of an image classification model. Specifically, for each class $y=y_{c}$, they first find regions in images with label $y_{c}$ that have maximum activation of the neuron $\theta_{i}$ as the surrogate explainees for $\theta_{i}$, and prompt LLMs such as ChatGPT to generate candidate explanations (words, short phrases) for the class label $y_{c}$. Then, they use the pretrained vision-language model CLIP (Radford et al., 2021) to match the candidate explanations with the surrogate explainees as the explanations for the neuron $\theta_{i}$. Recently, LLMs have also found applications in explaining recommender systems (Zhu et al., 2023). Specifically, (Yang et al., 2023b) found that LLMs can well interpret the latent space of sequential recommendation model after alignment, whereas (Lei et al., 2023) propose to align user tokens of LLMs with the learned user embeddings of small recommendation model to generate explanations of user preferences encoded in the embeddings. Recently,  (Schwettmann et al., 2024) propose a unified framework to explain all models where inputs and outputs can be converted to textual strings. Specifically, the explainer LLM is used as an agent to interact with the explainee model by iteratively creating inputs and observing outputs from the model, where the textual explanations are generated by viewing all the interactions as the context.
9.2. Self-Explanation of LLMs
@@
Due to the black-box nature of LLMs, it is promising to generate user-friendly explanations for the LLMs themselves, such that their operational mechanics and the predictions can be understood by human experts. Based on whether the LLM needs to be retrained to generate explanations for themselves, the self-explanation of LLM can be categorized into two classes: fine-tuning based approach and in-context based approach, which will be introduced in the following parts.
@@
Fine-tuning based approaches. Given sufficient exemplar explanations on the labels of the training data (e.g., in recommendation datasets such as the Amazon Review datasets (He et al., 2017) or the Yelp dataset (Zhou et al., 2020), users have provided explanations on why they have purchased certain items, which can be viewed as explanations for the ratings), LLMs can learn to generate explanations for their predictions as an auxiliary task through supervised learning. One exemplar method is P5 (Geng et al., 2022), which fine-tunes the pre-trained language model T5 (Raffel et al., 2020) on both the rating and explanation data to generate an explanation alongside the recommendations. Recently, several works have improved upon P5 (Cui et al., 2022; Zhu et al., 2024), which fine-tunes different LLMs such as GPT2, LLaMA, Vicuna, etc., and propose different prompt learning strategies (Li et al., 2023e) with generating explanation as the auxiliary task. With explanations introduced as additional supervision signals to fine-tune pretrained LLMs for recommendations, the performance can be improved with good explainability.
@@
In-context based approaches. In many applications, there is often a lack of sufficient exemplar explanations. However, the unique capability of modern LLMs to reason and provide answers through human-like prompts introduces the potential for in-context based explanations. Here, explanations for predictions are crafted solely based on the information within the prompt. A leading approach in this domain is the Chain-of-Thoughts (CoT) prompting (Wei et al., 2022), which provides few-shot examples (with or without explanations) in the prompt and asks the LLM to generate answers after reasoning step-by-step, where the intermediate reasoning steps that provide more context for generating the final answer can be viewed as explanations. However, CoT generates reasoning first and then based on which generates predictions, where the reasoning steps can influence prediction results (Lyu et al., 2023). If explanations are generated after the prediction, since the explanation is conditioned on the predicted label, it can provide a more faithful post-hoc explanation of why the model makes certain decisions (Lanham et al., 2023). The application of in-context based self-explanation of LLMs is broad. For example, (Huang et al., 2023c) explore generating zero-shot self-explanation of sentiment analysis with LLMs by directly asking them to generate explanations alongside the predictions. In addition, (Huang et al., 2023a) propose a chain-of-explanation strategy that aims to explain how LLMs can detect hate speech from the textual input. (Lu et al., 2022) find that CoT can generate well-supported explanations for question answering with scientific knowledge.
9.3. Challenges
9.3.1. Usability v.s. Reliability